Archive for the 'Repaired' Category

Light Shorted

Tuesday, October 13th, 2009

Paddock Pool LightToday I finally broke down and pulled the light from the deep end. It had thrown the GFI breaker last November and Paddock said it was not covered by warranty so I decided to take a look.
It seems that the rust proof housing had rusted through and the light filled with water. I am going to let it dry out and fill the holes in the can to see if it will work again.
It will be hard to replace since Paddock ran the wire through a maze of pipes back to the power panel.

SWG Installed

Thursday, May 28th, 2009

Finished installing the salt water chlorine generator. After refilling the pool, salt level down to 3400. SWG seems to be running fine. Set it to super chlorinate to see what happens.

New Chlorinator

Wednesday, May 27th, 2009

The salt water generator came in today. I mounted the controller to the wall and started draining the pool. I figured I will need to drain 25% of the water to get the salt level down to the new level for the new SWG.
I removed the worthless ozone generator. The cost to repair is much greater than the benefits.

Heatpump works

Tuesday, May 19th, 2009

Gillespie A/C called me and had me do a reset on the heatpump. This was basically tuning off the power to the heatpump and letting it sit, then restarting it. That seemed to fix it. It restarted with no errors. You would think it ran on Windows XP.

Heatpump Error

Monday, May 18th, 2009

Paddock came out and repaired the leaking pipe at the filter. Another $65.00 service charge.

When I restarted the pool pump after waiting about an hour for the glue to dry, the heatpump error code popped up saying the suction TH open. The heatpump refused to start. Gillespie A/C said that the temp sensor has failed and needs to be repaired. They will try and be out on Wednesday.

I called Pentair to ask if the heatpump could be replaced instead of constantly repairing the unit. After being on hold for about 15 minutes the service technician’s computer crashed and I was put back in the hold lineup. Another 15 minutes on hold. I was told that they would look into it and call me back Tuesday.

So, as of today, the leak is fixed but the chlorine generator (Ecco-Matic), Ozone generator (Ultra-Pure), Deep end light (Jazz Light), and heatpump (Pentair) have all failed.

Junk

Friday, May 8th, 2009

Paddock forgot to clean linesI blew out the main drain today, I had to hook up my air compressor to the drain filter and blow out the debris.

Since Paddock was sure it was not a warranty problem I decided to do it myself.
I found a chunk of plaster that was stuck inside the drain pipe. I wonder how much more plaster is stuck in the pipes.

The Ecco-Matic chlorinator still does not work and I have to come up with around $1200 to replace it with another brand that will last longer than 2 years. For now I have to use a floating chlorine dispenser.

I still need to pull the light in the deep end as it is still shorted.

I noticed that the pipe exiting the filter is leaking at an elbow. I wonder if Paddock considers bad glue joints warranty? I will have to call and ask.

Called Paddock and spoke with Debbie. The soonest available time is on May 18, 2009, and a cost of $65 for the trip.
I asked her about the light and she informed me that it has a warranty for 3 years. Unfortunately that expired in March.

Heat pump Repair

Thursday, March 5th, 2009

Gillespie A/C called and quickly came out to look at the unit. They determined that a capacitor had failed and replaced it. I am leaving the unit to run for a few days to see if it will continue to operate. Gillespie A/C was very helpful and their response was quick.

Air Release Repair

Wednesday, June 4th, 2008

Paddock came out and replaced the air release valve on top of the filter. They also tightened up the fittings that were leaking.

Heater Repair

Wednesday, April 23rd, 2008

Gillespie A/C came out and added freon to the system again. Will have to wait and see if that worked.

It has had service on:
06-18-06
07-03-06
08-14-06
12-09-06
12-15-06
04-12-08
04-23-08

2007 the pool was being repaired so the heater was never used.

New Heatpump Board

Tuesday, April 15th, 2008

Gillespie A/C came out and replaced the auto set board in the heatpump again. This I believe is the 4th board they have replaced.

Fixed PCC

Tuesday, August 21st, 2007

Fixed the PCC cleaning jet. A large piece of plaster had wedged itself under it keeping it from retracting. Blew out the line and it seems to be working again.

Electricians

Monday, June 25th, 2007

The electricians came out and rewired the pump.

Deck Touched Up

Friday, April 27th, 2007

The pool deck was touched up and looks much better.

Deck Cleaned

Thursday, April 26th, 2007

DSC00141The pool deck was cleaned today. Some of the plaster dust in the pool deck was removed and most of the rest are scrapes where the paint had been scratched off the deck. There is a quarter size check of deck missing in one spot.

Paddock Cleaned Deck

Thursday, April 19th, 2007

Paddock came by and cleaned the pool deck. It had white plaster though-out from when they plastered the pool. I will have to see if it cleaned up in the morning.

Replace Tiles

Wednesday, April 11th, 2007

Prado and Sons replaced the broken tiles and refilled the pool.

Fiber Optics

Thursday, March 1st, 2007

DSC00076Fiber Creations of Arizona came out today and glued the fiber optics back into the waterfall, again. This time he said they would also silicone around it to keep it from falling out, again. He knew nothing about having to replace it as was discussed in previous meetings.

Burn

Friday, February 9th, 2007

Prado and Sons came out later and power sanded the plaster. They then used a propane torch to try and pull out all moisture from the step.
After they left I called Debbie Moles and left a message. The footprint was still there, not as noticeable, but still there.

Heater repair

Friday, December 15th, 2006

Gillespie came out and changed the low pressure cutout saftey. He replaced the 90# unit with a 50# one.

Heatpump Repair

Saturday, December 9th, 2006

Gillespie A/C and Heating came out today and repaired the heat-pump. They added more refrigerant and replaced a pressure switch.

Heatpump Repair

Monday, August 14th, 2006

Ken with Westly Air-conditioning came out and repaired the heat-pump. He said a micro switch was malfunctioning.

Heatpump was Repaired

Monday, July 3rd, 2006

Heatpump was repaired. The refrigerant level was too high. They also replaced the control board with a new updated one. He also found the plug to the changeover valve had been disconnected, so it would never have changed over to cooling.

Re-grouted

Monday, June 19th, 2006

Waterline tile was re-grouted.
I called Kimberly and left a message about the heat pump failing.
She called back and said that the subcontractor will call me and schedule a time to look at it.

Another Pump

Tuesday, June 13th, 2006

The service tech came out and replaced the wok pot pump. He also blew out the main drain again, removing several large chunks of plaster from inside the line.
The fiber optics in the waterfall was glued back in place again.

I rebuilt the electrical panel

Thursday, May 25th, 2006

DSC00309I rebuilt the electrical panel. Most of the lights could not be controlled manually and the switches for the water feature pumps were all over the place. I moved all the confusing mess of switches into their own box. Now they can be found without a map. I added several switches for the lights and color wheels so they could be switched manually if the remote is lost or dies.
DSC00312The pump for the wok pots still shuts off and on whenever it wants. Either the pump is overheating or the switch is failing.

Fiber Optics

Monday, May 1st, 2006

DSC00304The fiber optic people came out and glued the fiber optics back into the waterfall.
I changed out the rock in the wok pots from river rock to a smaller beach pebble. It now matches the look of the pots in Paddocks showroom wok pots. I cut down the pipes in the wok pots below the water line. It now flows more water while keeping a bubbling type effect.

Repairs

Friday, April 28th, 2006

Electrician came out and set the fiber optics on the remote. Also seperated the color wheels.
Plastic pipe scrap removed from under wok pots.
Another Jandy valve top was replaced.

Pool Completed

Wednesday, April 26th, 2006

I met with Edward Ghiz, Debbie Moles, Bob and Chris from Poco Verde here. I tried to explain how the waterfall was supposed to look like. The one in their showroom has the waterfall set into a cantilever edge, hiding the sheer and lights from view. We looked over the pool and I explained the problems left to repair. Most all of the items had already been taken care of this week. All that is left is getting the fiber optics onto the remote, separating the color wheels for all the lights on the remote, fix the fiber optics falling out of the waterfall, and removing the white spacers on the wok pots. There is air getting into the filter system, causing it to loose prime, and he said they would fix that. One of the Jandy valves is malfunctioning also.
I showed him where the stucco on the face of the planter looked patched to me. Poco Verde offered to fix it but I refused, as they do not have a good track record with me on their repairs. Mr. Ghiz offered to have another contractor repair it and bill them the cost of the repair.
Mr. Ghiz said that the pool was complete and the rest is warranty work that they should have completed within a week.

Deck

Tuesday, April 25th, 2006

Pool deck people came out and re-sprayed the deck.

Acid Wash

Friday, April 21st, 2006

DSC00272Prado and Sons came out and acid washed the plaster.
They said they cleaned it as best they could.
Refilling the pool with water.

A List

Wednesday, April 19th, 2006

DSC00231Prado and Sons came out to look over the perma-dirt stains in the plaster. He said they would need to drain the pool to acid wash it. I told him that we had thought that they were going to drain it last week or the week before. This is about the third time they have been by to look at this. He said it was the first he had heard about it. He would get back with Paddock on what to do next.
DSC00239Paddock came out to resurface the pool deck and repair the chips. He was under the impression that the pool had been completed. They power washed the deck and left.
Ron with Paddock came out to check on the cleaning system. He was wondering if it had been drained yet. I told him that Prado and Sons had shut down the cleaning system earlier today. He would get back with Paddock on what to do next.
DSC00248A couple of guys showed up to look at the tile. They said it was not due to their caulking but it looked as if the grout had run on the tile. They would see if they could clean it.
Poco Verde came out to touch up the paint on the waterfall.

Things that are yet to be completed:
1.Color wheels not able to change separately on main lighting.
2.Fiber optic controller not on remote and change color wheel separately.
3.Concrete riser for wok pot # 3 fiber optics not finished.
4.Gauge for waterfall pump broken.
5.Water leak at Jandy valve for one of the wok pots.
6.Stains and chips on deck.
7.Cleaning system plugged with construction material.
8.Fiber optic lines falling out of waterfall.
9.Wok pot water lines can be seen (pipes above nozzles)
10.Stains in plaster. (perma-dirt)

I called Edward Ghiz at 3:00pm and asked for him or a representative of Paddock Pools to come out and look at the pool. He stated that he had to leave for the Las Vegas office but would come out next week. He said he would call me back with a time to meet.

Complaint filed with B.B.B.
On March 31, 2005 I contracted for a custom pool with Paddock Pools. It is now April 19, 2006 and they have yet to complete the pool.
I was forced to file a complaint with the R.O.C. on December 21, 2005 due to poor construction and they promised if I rescinded the complaint the pool would be completed shortly after February 15, 2006. I pulled the complaint and most of the items had been repaired.
Currently there are several items yet to finish.
( SEE ABOVE )
Other problems keep arising as construction continues.

These were promised to be complete by April 19, 2006 by Edward Ghiz, Sr Vice President, Construction Division.
I have a custom waterfall that has had to be rebuilt numerous times due to poor construction and was just refinished again this week.
I feel the time it has taken to build this pool is way beyond a reasonable amount. I have made several concessions to try and expedite the work.

I would just like them to finish the work as promised with the quality and workmanship expected from Paddock Pools.

Waterfall Repair

Tuesday, April 18th, 2006

DSC00216Poco Verde came out and repaired the waterfall face. They chipped away the falling stucco and patched the areas. They also replaced the rock into the wok pots.
Paddock called and the deck is scheduled for repairs on Wednesday.

PCC and Waterfall

Monday, April 17th, 2006

Ron came out and worked on the cleaning system.
Red Rock came out to see if they could find the sewer line to be able to drain the pool.
Poco Verde stopped by to touch up the paint on the patches in the waterfall. I showed him where the stucco is releasing from the wall. He said he would get in touch with Scott Dilliam and let him know he would be out to repair it Tuesday or Wednesday.

Week 54:

Thursday, April 6th, 2006

DSC00188The tile cleaner came by and tried to clean the tile. He said that he was able to get the streaking off but could not clean the grout stains in the tile itself. He did not want to try and clean the tile with harsh chemicals and destroy the tiles. He will check with Terry on what to do next.

Oops

Friday, March 17th, 2006

The electrician repaired the incorrectly wired pumps. He also said that the pool lights were also wired wrong and he fixed that also. He said they also installed the wrong lights in the pool. The color wheels can not be synced together, only one light has the auto sync feature the other light is a standard light. While checking his paperwork he also noted that they had installed Jazz lights, his paper calls for F/X lights. The F/X is a lower output lamp, and on a pool this size why lower output lamp? My contract states Jazz lights. So he needs to call and have a diver come out and replace the lights with the correct ones. They will not be able to operate the lights independently of each other the way they are wired and he doesn’t know if they will be able to sync colors with the fiber optics. According to Fiberstars, http://www.fiberstars.com/, all their lighting should be able to sync with each other assuming you get the right ones.
Jesse with Poco Verde called and said they are not able to seal the wok pots till the fiber optics are done in the pots.
Paddock came by to touch up the chips, stains, and cracks. He looked over the deck and said that he could do the repairs but since the construction was not finished, as he was told, he wanted to wait. He said that he would need to pressure wash and repaint the whole deck. I agreed that we need to wait because they still need to move rock and dirt across the deck.
Prado and Sons Plaster called and will be out on Tuesday to look over the stains in the plaster.
Water level down over 1/2″ since I turned off auto leveler. Pumps did not run all last night because the U/V system was installed.
Called Debbie and left another message.
Debbie Moles called and I gave her the list of things yet to be done. She was under the impression it was complete. She is going to schedule the electricians to figure out the lighting, fiber optic people, Poco Verde to seal the wok pots, Allen Tile to clean up the tile, the deck cleanup, and the plasterer to fix the stains. I am going to give them till the end of March to finish the project, at that time I will have no further contact with them.

Week 49

Thursday, March 2nd, 2006

Repaired hose bib next to pool. Debbie Moles called to check on hose bib. It was the hold up for the past 2 weeks. I explained to her my frustration again. In 2 weeks, just some paint and a hose bib.

Day 30 of contract – Final Day

Wednesday, February 15th, 2006

Fiber optics in pot #3 replaced. Optics could not be run in conduit. A cut into the concrete at the wok pot base so that the optics line could be rerun into the wok pot. This will place the fiber optics inside the concrete base. If any service is ever needed to the cable, the base would need to be torn up to get to the cable. Debbie Moles called and left a message, she will call again on Thursday.

Search

Recent Posts

Archives